OpenStudy (anonymous):

how to solve log4^16

jimthompson5910 (jim_thompson5910):

4 is the base right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

how would you write 16 as a power of 4?

OpenStudy (anonymous):

oh

jimthompson5910 (jim_thompson5910):

ie what is x in 4^x = 16

OpenStudy (anonymous):

4^2

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so log(4, 16) = log(4, 4^2) log(4, 16) = 2*log(4, 4) log(4, 16) = 2*1 log(4, 16) = 2

jimthompson5910 (jim_thompson5910):

or you can use the change of base formula to get log(4, 16) = log(16)/log(4) log(4, 16) = 2

OpenStudy (anonymous):

ok thnaks i got it

jimthompson5910 (jim_thompson5910):

you're welcome